# KGGen: Extracting Knowledge Graphs from Plain Text with Language Models

### v1

@@ Abstract
Recent interest in building foundation models for KGs has highlighted a fundamental challenge: knowledge-graph data is relatively scarce. The best-known KGs are primarily human-labeled, created by pattern-matching, or extracted using early NLP techniques. While human-generated KGs are in short supply, automatically extracted KGs are of questionable quality. We present a solution to this data scarcity problem in the form of a text-to-KG generator (KGGen), a package that uses language models to create high-quality graphs from plaintext. Unlike other KG extractors, KGGen clusters related entities to reduce sparsity in extracted KGs. KGGen is available as a Python library (pip install kg-gen), making it accessible to everyone. Along with KGGen, we release the first benchmark, Measure of of Information in Nodes and Edges (MINE), that tests an extractor’s ability to produce a useful KG from plain text. We benchmark our new tool against existing extractors and demonstrate far superior performance.
footnotetext: Code for this project can be found at https://github.com/stair-lab/kg-gen
@@ 1 Introduction
Knowledge graph (KG) applications and Graph Retrieval-Augmented Generation (RAG) systems are increasingly bottlenecked by the scarcity and incompleteness of available KGs. KGs consist of a set of subject-predicate-object triples, and have become a fundamental data structure for information retrieval (Schneider, 1973). Most real-world KGs, including Wikidata (contributors, 2024), DBpedia (Lehmann et al., 2015), and YAGO (Suchanek et al., 2007), are far from complete, with many missing relations between entities (Shenoy et al., 2021). The lack of domain-specific and verified graph data poses a serious challenge for downstream tasks such as KG embeddings, graph RAG, and synthetic graph training data.
Embedding algorithms such as TransE (Bordes et al., 2013) rely on abundant relational data to learn high-quality KG representations. In particular, TransE represents relationships as vector translations between entity embeddings and has demonstrated strong performance in link prediction when trained on large KGs (e.g., 1M entities and 17m training samples). However, if the KG is sparse or incomplete, embedding models struggle – they cannot learn or infer missing links effectively, degrading performance on knowledge completion and reasoning tasks (Pujara et al., 2017; Pote, 2024).
Consider retrieval-augmented generation (RAG) with a language model (LM) – this requires a rich external knowledge source to ground its responses. For instance, GraphRAG integrates a KG into the RAG pipeline (Edge et al., 2024). In GraphRAG, a language model (LM) like GPT-4o is used to extract a KG from a text corpus automatically, and this graph is used for retrieval and reasoning. This structured, graph-based augmentation has been shown to improve multi-hop reasoning and synthesis of information across documents (Larson & Truitt, 2024). By traversing relationships in the constructed graph, GraphRAG can “connect the dots” between disparate pieces of information, outperforming baseline RAG that relies only on semantic search over text. However, GraphRAG’s performance ultimately depends on the quality of the extracted graph (Zhang et al., 2024). In practice, automatically constructed graphs can be noisy and incomplete – some false nodes and edges may be introduced and some important ones omitted, which can hinder downstream reasoning (Thakur, 2024).
An emerging line of work that builds on graph-based RAG trains neural networks on KG retrieval. For example, GFM-RAG (Graph Foundation Model for RAG) (Luo et al., 2025) trains a dedicated graph neural network on an extensive collection of KGs, encompassing 60 graphs with over 14 million triples to serve as a foundation model for graph-based retrieval. By learning from diverse KGs, GFM-RAG’s retriever can generalize to unseen graphs and better handle the noise/incompleteness in automatically extracted KGs. These efforts underscore the importance of having dense, well-connected KGs to feed into RAG systems.
In this work, we propose KGGen (Text-to-Knowledge-Graph), a package that leverages LMs and a clustering algorithm to extract high-quality, dense KGs from text. KGGen addresses knowledge scarcity by enabling the automatic construction of KGs from any textual source rather than being limited to pre-existing databases like Wikipedia. The package uses an LM-based extractor to read unstructured text and predict subject-predicate-object triples to capture entities and relations. KGGen then applies an iterative LM-based clustering to refine the raw graph. Inspired by crowd-sourcing strategies for entity resolution (Wang et al., 2012), the clustering stage has an LM examine the set of extracted nodes and edges to identify which ones refer to the same underlying entities or concepts. Variations in tense, plurality, stemming, or capitalization are normalized in this process - e.g., “labors” might be clustered with “labor” and “New York City” with “NYC.” The resulting KG has far less redundancy and is densely interlinked, making it suitable for downstream use.
In addition to KGGen, we provide the first benchmark to measure text-to-knowledge-graph extraction. Our benchmark feeds $100$ Wikipedia-length articles into a KG extractor, then uses RAG to answer questions about the articles. On our benchmark, KGGen outperforms leading existing text-to-KG extractors by $18\%$. KGGen paves the way for a data-rich future when training next-generation KG foundation models and RAG systems.
@@
    We develop the first-ever benchmark for text-to-KG extractors, allowing for a fair comparison of existing methods.
@@
    We show that KGGen outperforms existing extraction methods by $18\%$ on this benchmark, exhibiting its potential to produce functional KGs using LMs.
@@ 2 Existing Methods
Before describing KGGen, we explain the two leading existing methods for extracting KGs from plain text, which will serve as a basis for comparison throughout the rest of this paper.
@@ 2.1 OpenIE
Open Information Extraction (OpenIE) was implemented by Stanford CoreNLP based on Angeli et al. (2015). It first generates a “dependency parse” for each sentence using the Stanford CoreNLP pipeline. A learned classifier then traverses each edge in the dependency parse, deciding whether to create (Yield), continue (Recurse), or stop processing a clause. These decisions split complex sentences into shorter, self-contained clauses. From these clauses, the system produces (subject, relation, object) tuples, each accompanied by a confidence score. Because OpenIE does not require its input text to have a specific structure, OpenIE can handle text in any format.
@@ 2.2 GraphRAG
Microsoft developed GraphRAG, which integrated graph-based knowledge retrieval with language models (LMs) Larson & Truitt (2024). As a first step, GraphRAG provides functionality for generating KGs from plain text to use as its database. In this process, GraphRAG creates a graph by prompting LMs to extract node-entities and relationships between these entities to serve as edges between the nodes. Throughout this extraction, few-shot prompting provides the LM with examples of “good” extractions. GraphRAG aggregates well-connected nodes into “communities” and generates a summary for each community to remove redundancy. The final graph consists of the communities as nodes and sentences summarizing their relationships as edges.
3 KGGen: KGs From Plain Text
Unlike most previous methods of LLM-based KG extraction, we rely on a multi-stage approach involving an LLM (in our case, GPT-4o) to (1) extract entity and relations from each source text, (2) aggregate graphs across sources and (3) iteratively cluster entities and relations. We implement these stages in a modular fashion via a new kg-gen Python toolkit consisting of a ‘generate‘ module for extraction, an ‘aggregate‘ module for source consolidation, and a ‘cluster‘ module for dynamic entity resolution. We use the DSPy framework throughout these stages to define signatures that ensure that LLM responses are consistent JSON-formatted outputs. In our case, we use GPT-4o, although the implementation may be used with any model supported by DSPy.
We impose strong constraints on the LLM via prompting to reduce the likelihood of semantically dissimilar duplicate entities. We introduce multiple passes through our extracted edges and relations to cluster similar entities and consolidate the number of edge types. Consolidation and clustering prevent the formation of sparse KGs, which may produce meaningless KG embeddings under standard algorithms such as TransE.
@@ 3 KGGen: KGs From Plain Text
3.1 Entity and Relation Extraction (‘generate‘)
The first stage takes unstructured text as input and produces an initial knowledge graph as extracted triples. We invoke the GPT-4o model for each input text through a DSPy signature that instructs the model to output detected entities in a structured format. Then, we invoke a second LLM call through DSPy that instructs the model to output the subject-predicate-object relations, given the set of entities and source text. We find this 2-step approach works better to ensure consistency between entities.
3.2 Aggregation (‘aggregate‘)
@@
3.3 Entity and Edge Clustering (‘cluster‘)
After extraction and aggregation, we typically have a raw graph containing duplicate or synonymous entities and possibly redundant edges. The clustering stage is a key innovation in our KG extraction methodology that aims to merge nodes and edges representing the same real-world entity or concept. We take an iterative LLM-based approach to clustering, inspired by how a group of humans might gradually agree on consolidating terms. Rather than attempting to solve the entire clustering in one shot (which is intractable for an extensive list of entities), KGGen performs a sequential series of clustering operations for entities:
@@ 3.3 Entity and Edge Clustering (‘cluster‘)
    The entire entities list is passed in context to the LLM, and it attempts to extract a single cluster. An optional cluster-instruction string may be passed to decide how to cluster. The default instructions account for close synonyms and differences in tense and plurality.
@@
    Validate the single cluster using an LLM-as-a-Judge call with a binary response. If it passes, then add the cluster and remove the cluster entities from the entities list.
@@
    Assign a label to the cluster that most closely captures the shared meaning of entities in the cluster.
@@
    Repeat steps 1–3 until $n$ loops happen without a successful cluster extraction.
  5. 5.
    Remaining entities are checked batch-by-batch, with batch size $b$, for whether they should be added to an existing cluster.
  6. 6.
    For each new addition to a cluster, validate the cluster once more using an LLM-as-a-Judge call with a binary response.
  7. 7.
    Repeat steps 5–6 until there are no remaining entities to check.
The same operations are performed on edges, albeit with slightly modified prompts.
The clustering process allows us to create dense KGs that admit meaningful embeddings. To give a real example of the usefulness of our process, in one of our raw KGs, we found the entities “vulnerabilities”, “vulnerable”, and “weaknesses”. Although these are different words, they have similar meanings and should be viewed as equivalent in our KG.
@@
4 A Benchmark for Extraction Performance
@@
Although a handful of existing methods attempt to extract KGs from plain text, it is difficult to measure progress in the field due to the lack of existing benchmarks. To remedy this, we produce the Measure of Information in Nodes and Edges (MINE), the first benchmark that measures a knowledge-graph extractor’s ability to capture and distill a body of text into a KG.
4.1 MINE Description
MINE involves generating KGs for 100 articles, each representing a distinct source of textual data. Each article is approximately 1,000 words long and is generated by an LLM based on a diverse list of 100 topics that range from history and art to science, ethics, and psychology. To evaluate the quality of the generated KGs, we develop a metric to assess how effectively they capture critical information from the articles.
We extract 15 facts–here defined as statements present in the plain text article–from each article by providing an LLM with the article and the extraction prompt found in Appendix C. We manually verify that the 15 facts are accurate and contained in the article. MINE assesses how well a text-to-KG extractor captures the information present in the text by determining whether these 15 facts are captured by the KG generated from the article.
For each article, KGs are generated using using the plain-text-to-KG method being benchmarked. The nodes of the resulting KGs are then vectorized using the all-MiniLM-L6-v2 model from SentenceTransformers, enabling us to use cosine similarity to assess semantic closeness between the short sentence information and the nodes in the graph.
For each KG generation method, the KG for each article is queried for each of the 15 facts from that article. We do this by determining the top-k nodes most semantically similar to each fact. Next, we determine all the nodes within two relations of one of the top k-nodes. Finally, we return all these nodes along with their relations as the result of the query. This result is subsequently evaluated using an LLM, provided it is queried for and a specific prompt to produce a binary output: 1 if the fact could be inferred from only the information in the queried nodes and relations, and zero otherwise. The prompt can be found in Appendix C.
The final MINE score of each KG generator on a given article was calculated as the percentage of 1s across all 15 evaluations. This systematic approach objectively compares the methods based on their ability to capture and retrieve information from the articles accurately.
This evaluation process is illustrated in Figure 2.
Figure 2: Evaluation process used in MINE

### v2

@@ Abstract
Recent interest in building foundation models for knowledge graphs has highlighted a fundamental challenge: knowledge graph data is scarce. The best-known knowledge graphs are primarily human-labeled, created by pattern-matching, or extracted using early NLP techniques. While human-generated knowledge graphs are in short supply, automatically extracted ones are of questionable quality. We present KGGen, a novel text-to-knowledge-graph generator that uses language models to extract high-quality graphs from plain text with a novel entity resolution approach that clusters related entities, significantly reducing the sparsity problem that plagues existing extractors. Unlike other KG generators, KGGen clusters and de-duplicates related entities to reduce sparsity in extracted KGs. Along with KGGen, we release Measure of Information in Nodes and Edges (MINE), the first benchmark to test an extractor’s ability to produce a useful KG from plain text. We benchmark our new tool against leading existing generators such as Microsoft’s GraphRAG; we achieve comparable retrieval accuracy on the generated graphs and better information retention. Moreover, our graphs exhibit more concise and generalizable entities and relations. Our code is open-sourced at https://github.com/stair-lab/kg-gen/.
Belinda Mo∗1, Kyssen Yu∗2, Joshua KazdanEqual Contribution. 1, Proud Mpala1, Lisa Yu2, Chris Cundy3,
Charilaos Kanatsoulis1, Sanmi Koyejo1
1Stanford University   2University of Toronto   3 FAR AI
@@ 1 Introduction
Knowledge graph (KG) applications and graph-based Retrieval-Augmented Generation (RAG) systems are increasingly bottlenecked by the scarcity and incompleteness of available KGs. KGs consist of a set of subject-predicate-object triples, and have become a fundamental data structure for information retrieval (Schneider, 1973). Most real-world KGs, including Wikidata (contributors, 2024), DBpedia (Lehmann et al., 2015), and YAGO (Suchanek et al., 2007), are far from complete, with many missing relations between entities (Shenoy et al., 2021). The lack of domain-specific and verified graph data poses a serious challenge for downstream tasks such as KG embeddings, graph RAG, and synthetic graph training data.
Several synthetic plain-text-to-KG extractors have been proposed to address KG scarcity, most prominently including OpenIE (Angeli et al., 2015) and Microsoft’s GraphRAG (Larson and Truitt, 2024). Both OpenIE and GraphRAG extract entities and relations directly from text, but they lack effective mechanisms for entity resolution and relation normalization. This leads to graphs with nearly as many unique relation types as edges, resulting in sparse, disconnected knowledge representations that limit their utility for downstream tasks. To solve this problem, we propose KGGen, a text-to-knowledge-graph generator that leverages language models (LMs) and an algorithm for entity and edge resolution to extract high-quality, dense KGs from text. First, KGGen uses an LM-based extractor to read unstructured text and predict subject-predicate-object triples to capture entities and relations; after extracting the triples, it applies a novel, iterative clustering algorithm to refine the raw graph. Inspired by crowd-sourcing strategies for entity resolution (Wang et al., 2012), KGGen identifies nodes that refer to the same underlying entities, and consolidates edges that have equivalent meanings.
The nascent field of plain-text-to-knowledge graph extraction currently lacks benchmarks to measure the fidelity of KG generation from text. To close this gap, we provide two new benchmarks: the first captures information retention from short texts; the second, based on WikiQA, measures knowledge retrieval capabilities for graphs generated from multi-million token, web-based knowledge databases. On these benchmarks, KGGen performs comparably to GraphRAG. However, KGGen exhibits far better scaling properties with respect to information compression and graph sparsity as the plain-text database length increases.
@@
    We develop benchmarks to drive improvements in plain-text-to-knowledge-graph extraction, and measure KGGen’s performance on these benchmarks.
@@
    We show that KGGen exhibits improved scaling with respect to the size of the text source relative to past methods.
2 Related Work
Interest in automated methods to produce structured text to store ontologies dates back to at least 2001 when large volumes of plain text began to flood the fledgling internet (Maedche and Staab, 2001). KG extraction from unstructured text has seen significant advances through rule-based and LM-powered approaches in the last 15 years. Early work (Suchanek et al., 2007) used hard-coded rules to develop YAGO, a KG extracted from Wikipedia containing over five million facts, and rules-based extraction still has appeal for those producing KGs in multi-modal domains today (Norabid and Fauzi, 2022; Oramas et al., 2015). With the development of modern natural language processing, hard-coded rules generally ceded to more advanced approaches based on neural networks. For instance, OpenIE (Angeli et al., 2015) provides a two-tiered extraction system: first, self-contained clauses are identified by a classifier; then, (Angeli et al., 2015) run natural logic inference to extract the most representative entities and relations from the identified clauses. Stanford KBP (Angeli et al., 2013) presents another seminal early approach to using deep networks for entity extraction.
As early as 2015, some hypothesized that extracting KGs would go hand-in-hand with developing better language models (Domeniconi et al., 2015). More recently, evidence has emerged that transformer-based architectures can identify complex relationships between entities, leading to a wave of transformer-based KG extraction techniques, which range from fully automatic (Qiao et al., 2022; Arsenyan et al., 2023; Zhang and Soh, 2024) to human-assisted (Kommineni et al., 2024). Our contribution to the extraction literature is to build KGs conducive to embedding algorithms such as TransE and TransR (Bordes et al., 2013; Lin et al., 2015). We observed that when one extracts KGs from plaintext, the nodes and relations are often so specific that they are unique. This causes the estimation of embeddings to be under-specified. We develop a method for automatic KG extraction from plain text that clusters similar nodes and edges to prevent this under-specification. This leads to a KG with better connectivity and more functional nodes and edges.
Evaluating the quality of knowledge graphs is important to ensure usefulness and reliability in downstream applications. Early evaluation methods focused primarily on directly assessing aspects such as completeness and connectivity or using rule-based statistical methods, while recent approaches emphasize usability in downstream applications and incorporation of semantic coherence(Xue and Zou, 2023).
In the late 2000s, research focused on assessing the correctness and consistency of KGs. The evaluations relied on expert annotations by selecting random facts from the generated KG and then calculating the accuracy of those facts. (Suchanek et al., 2007) This proved to be laborious and prone to errors. This led to accuracy approximation methods like KGEval (Ojha and Talukdar, 2017) and Two-State Weight Clustering Sampling(TWCS) (Gao et al., 2018), which employed sampling methods with statistical guarantees as well as use less annotation labor. As the KGs became larger and more diverse, particularly with the rise of automated extraction techniques from web data, this generated more pressure on annotators, leading to methods like Monte-Carlo search being used for the interactive annotation of triples (Qi et al., 2022). Furthermore, because accuracy alone did not fully capture the complexity of the knowledge graph, more evaluation metrics like completeness were used to characterize the quality of knowledge graphs. (Issa et al., 2021).
Two notable approaches to KG evaluation are the LP-Measure and the triple trustworthiness measurement (KGTtm) model. LP-Measure assesses the quality of a KG through link prediction tasks, eliminating the need for human labor or a gold standard (Zhu et al., 2023). This method evaluates KGs based on their consistency and redundancy by removing a portion of the graph and testing whether the removed triples can be recovered through link prediction tools. Empirical evidence suggests that LP-Measure can effectively distinguish between “good" and “bad" KGs. The KGTtm model, on the other hand, evaluates the coherence of triples within a knowledge graph Jia et al. (2019). Based on these evaluation methods, frameworks like Knowledge Graph Evaluation via Downstream Tasks(KGrEaT) and DiffQ(differential testing) emerged. KGrEaT provides a comprehensive assessment of KGs by evaluating their performance on downstream tasks such as classification, clustering, and recommendation (Heist et al., 2023) rather than focusing solely on correctness or completeness. In contrast, DiffQ uses embedding models to evaluate the KG’s quality and assign a DiffQ Score, resulting in improved KG quality assessment Tan et al. (2024).
@@ 3 Existing Methods
Before describing KGGen, we explain two popular existing methods for extracting KGs from plain text, which will serve as a basis for comparison throughout the rest of this paper.
@@ 3.1 OpenIE
Open Information Extraction (OpenIE) was implemented by Stanford CoreNLP based on Angeli et al. (2015). It first generates a “dependency parse" for each sentence using the Stanford CoreNLP pipeline. A trained classifier then traverses each edge in the dependency parse, deciding whether to create, continue, or stop processing a clause. These decisions split complex sentences into shorter, self-contained clauses. From these clauses, the system produces (subject, relation, object) tuples, each accompanied by a confidence score. Because OpenIE does not require its input text to have a specific structure, OpenIE can handle text in any format.
@@ 3.2 GraphRAG
GraphRAG, developed by Microsoft in 2024, integrates graph-based knowledge retrieval with language models (LMs) (Larson and Truitt, 2024). As a first step, GraphRAG provides functionality for generating KGs from plain text, which serve as its database for retrieval. In this process, GraphRAG creates a graph by prompting LMs to extract node-entities and relationships between these entities. Throughout this extraction, few-shot prompting provides the LM with examples of desireable extractions. GraphRAG aggregates well-connected nodes into communities and generates a summary for each community. The final graph consists of the nodes and their relationships along with communities their summaries.
4 KGGen: Knowledge Graphs From Plain Text
Unlike most previous methods of LLM-based KG extraction, we rely on a multi-stage approach (1) extract entity and relations from each source text using an LLM, (2) aggregate graphs across sources, and (3) iteratively resolve duplicate entities and edges using a hybrid of LLM and traditional informational retrieval methods.
We impose strong constraints on the LLM via prompting to prevent it from incorrectly grouping together entities or edges that are similar in meaning but not actually the same - for example, conflating “Type 1 diabetes” and “Type 2 diabetes,” “hypertension” and “stress,” or “MRI” and “CT scan.”. We introduce multiple passes through our extracted edges and edges to resolve similar entities and consolidate the number of edge types. Entity and edge resolution prevents the formation of sparse KGs, which may produce meaningless KG embeddings under standard algorithms such as TransE.
@@ 4 KGGen: Knowledge Graphs From Plain Text
4.1 Entity and Relation Extraction
The first stage takes unstructured text as input and produces an initial knowledge graph as extracted triples. We use Google’s Gemini 2.0 Flash as the language model to provide structured output via DSPy signatures. The first step takes in source text and extracts a list of entities. Given the source text and entities list, the second step outputs a list of subject-predicate-object relations. Each step corresponds to a DSPy signature that specifies instructions for the LLM to follow in its docstring. We find this 2-step approach works better to ensure consistency between entities.
4.2 Aggregation
@@
4.3 Entity and Edge Resolution
After extraction and aggregation, we typically have a raw graph containing duplicate or synonymous entities and possibly redundant edges. The resolution stage is a key innovation in our KG extraction methodology that merges nodes and edges representing the same real-world entity or concept. Our resolution process employs a two-stage approach combining embedding-based clustering with LLM-based de-duplication to efficiently handle large knowledge graphs. The approach is applied to both entity and edge items separately:
First, all items in the graph are clustered. We get the semantic embeddings of every item using S-BERT and cluster using k-means into clusters of 128 items.
@@ 4.3 Entity and Edge Resolution
    For each item in a cluster, we retrieve the top-k most semantically similar items, where k=16, using a fused BM25 and semantic embedding approach.
@@
    Then, the LLM is prompted to identify exact duplicates from this set, considering variations in tense, plurality, case, abbreviations, and shorthand forms.
@@
    For each set of duplicates, the LLM selects a canonical representative that best captures the shared meaning, similar to aliases that Wikidata uses. Cluster maps track which entities belong to which alias.
@@
    The item and its duplicates are removed from the cluster and steps 1-3 repeat until no items remain in the cluster.
This approach enables effective de-duplication even for very large knowledge graphs by processing semantic clusters in parallel. When processing our largest 20M chunk dataset, this method successfully consolidated entities like "Olympic Winter Games", "Winter Olympics", and "winter Olympic games" into a single canonical representation.
@@
5 Benchmarks for Extraction Performance
@@
Although a handful of existing methods attempt to extract KGs from plain text, it is difficult to measure progress on this task due to the lack of existing benchmarks. As a remedy, we produce the Measure of Information in Nodes and Edges (MINE), the first benchmark that directly measures a knowledge-graph extractor’s ability to capture and distill a body of text into a KG. MINE contains two tasks: the first assesses how well a KG extractor captures the information in short, page-length text; the second measures downstream RAG performance on multi-million token datasets. We call these tasks MINE-1 and MINE-2 respectively. MINE-1 ensures that a KG extractor accurately represents the source text, and MINE-2 gauges the practicality of the knowledge graphs for realistic applications.
5.1 MINE-1: Knowledge Retention
MINE-1 approximates the fraction of information a KG extractor is able to capture from an article without relying on downstream tasks, which can obscure whether performance gains stem from the KG extractor itself or from aspects of the extraction process.
MINE-1 consists of 100 articles, each accompanied by 15 facts that are known to be present in the article. The dataset has the following characteristics: articles have a mean length of 592 words (std. 85 words, range: 440-976 words) and cover diverse topics including Arts, Culture & Society (24 articles), Science (27 articles), Technology (19 articles), Psychology/Human Experience (18 articles), and History & Civilization (17 articles). Articles are generated by an LLM to ensure balanced coverage across these domains. For each article, MINE-1 generates a corresponding KG using the extractor being evaluated.
To assess the quality of these KGs, we extract 15 facts from each article using an LLM-based extraction prompt found in Appendix B. We manually verify that the 15 facts are accurate and contained in the article. To measure performance on MINE-1, the KG extractor first extracts a KG from each article. Then, MINE-1 contains a process to verify whether each fact can be recovered from the corresponding KG.
Figure 2: Evaluation process used in MINE-1
Verification occurs via a semantic query process: both the 15 facts and all KG nodes are embedded using the all-MiniLM-L6-v2 model from SentenceTransformers. For each fact, the verifier retrieves the top-k most semantically similar nodes in the KG, then expands the result to include all nodes within two relations of those top-k nodes, reflecting the fact that KG’s are often used for multi-hop reasoning tasks. The subgraph induced by these nodes is passed to an LLM, which is prompted to output a binary score: 1 if the fact can be inferred from the retrieved nodes and relations alone, and 0 otherwise. The prompt is detailed in Appendix B. The final MINE-1 score for a KG extractor is the percentage of the 15 facts scored as 1, averaged across all 100 articles. While LLM-based evaluation introduces potential biases, we validated its reliability by manually scoring 60 randomly selected fact-KG pairs and comparing them to LLM judgments, achieving 90.2% agreement and a correlation of 0.80. The full evaluation pipeline is illustrated in Figure 2.
5.2 MINE-2: KG-Assisted RAG Description
The RAG evaluation is based on the WikiQA dataset Yang et al. (2015), which contains 20,400 questions based on 1,995 Wikipedia articles. Using the method under evaluation, we construct a single KG that aggregates information from all articles referenced in WikiQA. For each question in the dataset, we retrieve the top 10 most relevant triples by embedding both the question and all KG triples using the all-MiniLM-L6-v2 model from SentenceTransformers. We then compute the cosine similarity between the question and each triple, alongside a relevance score produced by BM25. The final similarity score is obtained by combining BM25 relevance score and the cosine similarity score, weighted equally. The 10 triples with the highest combined scores are selected, and we further expand this set by appending 10 additional triples that fall within two hops of the nodes in the top 10 triples to enable multi-hop reasoning.
Since each relation is linked to a source text chunk during generation by KGGen and GraphRAG, we provide the full set of 20 retrieved triples, their associated text chunks, and the original question to an LM, which synthesizes an answer based on these inputs. The complete prompt used can be found in Appendix B. Finally, the LM responses are evaluated using LLM-as-a-Judge to determine whether they contain the correct answer to the question. The prompt used for this final verification step is also included in Appendix B. OpenIE is excluded from this comparison, as it cannot produce KGs that link relations to the original text chunks.